OpenStudy (anonymous):

Evaluate lim u-> infinity (tan^-1 u)

OpenStudy (anonymous):

that would be \[\frac{\pi}{2}\]

OpenStudy (anonymous):

\[\lim_{u \rightarrow \infty}( \tan ^{-1}u)\]

OpenStudy (anonymous):

can u plz show some work

OpenStudy (anonymous):

im not sure how 2 approach it

OpenStudy (anonymous):

i am not sure what "work" there is. range of arctangent is \[(-\frac{\pi}{2},\frac{\pi}{2})\] because that is what you restrict the domain of tangent to make it one to one

OpenStudy (anonymous):

as \[\lim_{x\rightarrow \frac{\pi}{2}^-} \tan(x)=\infty\] then you get your limit

OpenStudy (anonymous):

so teh answer is pi/2 to the left side

OpenStudy (anonymous):



OpenStudy (anonymous):

your best bet is to look at the graph (not my crappy picture) of \[y=\tan(x)\] and of \[y=\tan^{-1}(x)\] and the limit will be clear

OpenStudy (anonymous):

i see. so how come u wrote pi/2 from te hleft side ...= infinity